# Perseus: Removing Energy Bloat from Large Model Training

### v1

@@ Abstract
Training large AI models on numerous GPUs consumes a massive amount of energy. We observe that not all energy consumed during training directly contributes to end-to-end training throughput, and a significant portion can be removed without slowing down training, which we call energy bloat.
@@
In this work, we identify two independent sources of energy bloat in large model training, intrinsic and extrinsic, and propose Perseus, a unified optimization framework that mitigates both. Perseus obtains the “iteration time–energy” Pareto frontier of any large model training job using an efficient iterative graph cut-based algorithm and schedules energy consumption of its forward and backward computations across time to remove intrinsic and extrinsic energy bloat. Evaluation on large models like GPT-3 and Bloom shows that Perseus reduces energy consumption of large model training by up to 30%, enabling savings otherwise unobtainable before.
1 Introduction
As deep neural networks (DNNs) continue to grow in model and dataset size [29, 25], the energy consumption of large model training is increasing as well. For instance, training GPT-3 [11] reportedly consumed 1,287 megawatt-hour (MWh) in 2021 [48]. The recent surge in the development of generative AI (GenAI) and large language models (LLMs) shows no signs of deceleration in AI energy demands.
Optimizing the energy consumption of large model training is, therefore, crucial to minimizing the operational costs associated with AI, with additional benefits. For example, low-carbon electricity, which is preferred especially in the context of sustainability commitments [1, 4, 6, 7], is still very much a scarce and contended resource [50]. Reducing energy consumption allows more workloads to fit within the limited supply of low-carbon electricity [10]. Furthermore, reducing energy consumption without slowdown reduces power draw, either allowing more aggressive datacenter power oversubscription or conversely reducing power throttling for existing workloads [35, 54, 46].
@@ 1 Introduction
Despite many recent works on increasing the throughput of large model training [26, 41, 19, 43, 73, 34, 27], reining in the growing energy consumption remains an open challenge [55, 48]. In this paper, we aim to identify and remove energy bloat – the portion of energy consumption that can be removed without throughput loss – in large model training. We identify two independent sources of energy bloat: intrinsic and extrinsic, and propose a single optimization framework that minimizes both.
@@
Intrinsic energy bloat comes from computation imbalance when a large model is distributed across multiple GPUs with pipeline parallelism (§2.2). Balancing the amount of computation in each pipeline stage is an important problem for distributed execution planning [41, 19, 23, 73], but perfectly balancing every stage is not always possible because layers in a DNN are coarse-grained tensor operations with varying amounts of computation. When all stages do not have the same computation intensity, those not on the critical path of computation run needlessly fast – i.e., they consume energy that does not contribute to the overall training throughput. Such intrinsic energy bloat opens up the opportunity to precisely slow down each non-critical computation in the pipeline such that the length of the critical path does not change.
@@
Extrinsic energy bloat, in contrast, arises when multiple pipelines run in parallel in a synchronous fashion to scale out training to massive datasets, and one or more pipelines run slower than the others (§2.3). Root causes behind such slowdowns are varied, including power/thermal throttling [75, 47, 35, 54, 46], I/O bottlenecks in the storage/network [39, 71], hardware/software failures [61, 27], etc., and the likelihood of their presence increases with the scale and duration of training [22, 28, 64]. All pipelines running faster than the straggler pipeline are needlessly fast and wasting energy that does not affect the overall training throughput. Extrinsic energy bloat allows us to slow down entire pipelines (while still keeping their intrinsic energy bloat low) without delaying gradient synchronization.
@@
In this work, we propose Perseus that relies on a unified optimization framework to remove both intrinsic and extrinsic energy bloat during large model training (§3). At its core, Perseus efficiently pre-characterizes the entire “iteration time–energy” Pareto frontier, which enables it to minimize intrinsic energy bloat under normal operating conditions and to mitigate extrinsic energy bloat arising from straggler pipelines. Existing works fall short on both fronts. EnvPipe [13] is limited only to intrinsic bloat reduction with a point solution that leads to suboptimal energy reduction. Zeus [68], in contrast, ignores intrinsic bloat altogether as it only considers single-GPU training, which also leads it to identify suboptimal “iteration time–energy” frontiers at best.
@@
We prove that optimally solving our optimization problem is in fact NP-hard and propose an efficient algorithm that provides an exact solution for a relaxed problem (§4). At its core, our solution represents an iteration of large model training as a directed acyclic graph (DAG). Each node in the DAG represents either a forward or backward computation in a particular stage, and nodes are connected with edges that represent computation dependencies. Importantly, each node is annotated with its planned execution time and energy consumption, which we call the energy schedule. Perseus’s optimization algorithm starts from the energy schedule that consumes the least possible amount of energy, and iteratively speeds up the execution time of the schedule while incurring the least possible energy increase, until no more speed up is possible. This generates all energy schedules on the “iteration time–energy” Pareto frontier. Minimizing intrinsic and/or extrinsic energy bloat is then as simple as choosing the appropriate energy schedule from the Pareto frontier.
While solving the optimization problem depends on accurate fine-grained computation energy profiling, modern GPUs provide energy consumption information only in coarse-grained time intervals. Perseus introduces an online profiling technique that can accurately profile the energy consumption of fine-grained computations with low overhead (§5).
@@
Evaluation on a variety of large models including GPT-3 [11], BERT [15], T5 [51], Bloom [66] and a scaled-up version of Wide-ResNet [69] shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown (§6).
@@
  • We propose Perseus, an open-source unified optimization framework that aims to remove both types of energy bloat using an efficient graph cut-based algorithm.
  • We evaluate Perseus on a diverse set of large model workloads and show that it significantly reduces energy bloat.
2 Motivation
First, we provide necessary background regarding large model training (§2.1). Then, we introduce intrinsic (§2.2) and extrinsic (§2.3) energy bloat present in large model training.
2.1 Large Model Training
The rapid increase in DNN model size in recent years made data, tensor, and pipeline parallelism [26, 41, 43, 19, 73, 34] essential ingredients for large model training. Especially, pipeline parallelism partitions a large model into multiple stages and assigns each stage to different GPUs. The training batch is split into multiple microbatches, and their forward and backward computations are pipelined through those stages. Then, in order to scale out to more GPUs, such pipelines are replicated multiple times to perform data parallel training, where each pipeline computes a subset of the training batch and synchronizes gradients with one other at the end of each iteration. Due to the synchronization barrier in the end, the next iteration can only begin when all pipelines have finished.
2.2 Intrinsic Energy Bloat
Figure 1: One training iteration of GPT-3 1.3B with 4 pipeline stages (partitioned with minimum imbalance ratio) and 6 microbatches on NVIDIA A100 GPUs drawn to scale. For example, F5 and B5 in the S2 row denote Forward and Backward for the fifth microbatch on Stage 2. The critical path of computation is traced with a blue line. Colors are average power consumption. The blue background indicates that GPUs consume power when blocking on communication.
Model Size Imbalance Ratio
@@ 2.2 Intrinsic Energy Bloat
Wide-ResNet50 [69] 0.8B 1.23 1.46
Wide-ResNet101 [69] 1.5B 1.09 1.25
@@
Table 1: Imbalance ratio between the longest and the shortest stages for various models on NVIDIA A100 GPUs. 1.00 would mean perfect balance.
@@
We profile GPT-3 1.3B on NVIDIA A100 GPUs and visualize the timeline of one training iteration in Figure 0(a). In addition to the familiar bubbles found in any pipeline schedule [41], we observe gaps in between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage. Such gaps exist because the computation time of each pipeline stage is not perfectly balanced. Partitioning stages in a balanced manner is an important problem in distributed execution planning [26, 41, 19, 73], but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying size.
@@
To understand the amount of possible pipeline stage imbalance, we exhaustively enumerated all possible pipeline partitions [^1] For Transformer-based models, we partition at the granularity of Transformer layers. For Wide-ResNet, we partition at the granularity of bottleneck layers, which are three Convolution layers wrapped with a Skip connection. to find the one with the smallest imbalance ratio, defined as the ratio of the longest stage forward computation latency to the shortest. Table 1 lists the minimum imbalance ratio for various models, which shows that perfect balance is difficult to achieve. Appendix A explains the partitioning details of each model and where imbalance comes from.
@@
Given some amount of stage imbalance, as we have seen in Figure 0(a), not all forward and backward computations are on the critical path of computation. This means that non-critical computations running at their maximum speed is not contributing to faster iteration time, and thus simply wasting energy. We call this intrinsic energy bloat, which can be reduced by precisely slowing down each non-critical computation so that they do not affect the critical path (Figure 0(b)). Although seemingly simple, we prove that this problem is NP-hard by reduction from the 0/1 Knapsack problem (Appendix B). Visualizations for other models are in Appendix C.
2.3 Extrinsic Energy Bloat
@@
During large model training, numerous replicas of the same pipeline run in a data parallel fashion in order to train on more data with more GPUs, and every pipeline must synchronize gradients at the end of each iteration before moving on to the next. When one pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes. Figure 1(a) illustrates this situation. Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time are wasteful. We call this extrinsic energy bloat, because unlike intrinsic energy bloat, the cause of energy bloat is extrinsic to the pipeline.
Perseus finds the energy-optimal iteration time for the non-straggler pipeline, again by precisely slowing down computations in the pipeline (Figure 1(b)). However, unlike other problem settings that do not consider energy consumption, we show in Section 3.1 that fully utilizing all the slack time created by the straggler may not be energy-optimal! Perseus can automatically make such decisions and implements the energy-optimal slowdown.
@@
Figure 2: Toy example illustrating one straggler pipeline running together with a non-straggler pipeline using data parallelism. The non-straggler wastes energy by running as fast as possible. Perseus can precisely slow down non-straggler pipelines to reduce such extrinsic energy bloat.
Stragglers arise from numerous sources, with many introducing sizable energy bloat according to recent reports. Thermal or power throttling in a datacenter may lower the GPU’s frequency to prevent hardware or the power supply unit from being damaged, resulting in 10–50% slowdown [75, 47, 35, 54, 46]. I/O bottlenecks in the storage or network, especially during data loading for image, video, or recommendation models, can be longer than GPU computation by up to 4×4$\times$4 × [39, 71], thereby acting like a persistent (virtual) straggler pipeline. Recent failure-resilient large model training frameworks [61, 27] deploy heterogeneous pipelines with different number of stages or amount of compute after failures, introducing non-uniform iteration times across pipelines until recovery. As the scale of large model training jobs increase, the probability of the job encountering one or more causes of stragglers increases [22, 28, 64].
Stragglers due to events like thermal/power throttling, consistent I/O bottlenecks, and fault-tolerant planning can persist across many iterations [46, 39, 27]. A training system can identify such stragglers and calculate the extent of slowdown by comparing iteration times across data-parallel pipelines. In this work, we assume that such knowledge is available and focus on how to plan time and energy consumption across time and allow quick adaptation given such information.
3 Perseus Overview
@@
3.1 Optimization Approach
Figure 3: Three cases of where the straggler pipeline’s iteration time $T^{\prime}$ can be. The first case is when there are no stragglers. $T_{\min}$ and $T^{*}$ are the shortest and longest iteration time on the “iteration time–energy” Pareto frontier. The black dot is the default mode of execution where all computations run at the maximum speed. The green dot is the energy-optimal iteration time of the non-straggler pipeline.

### v2

# Reducing Energy Bloat in Large Model Training

@@ Abstract.
Training large AI models on numerous GPUs consumes a massive amount of energy, making power delivery one of the largest limiting factors in building and operating datacenters for AI workloads. However, we observe that not all energy consumed during training directly contributes to end-to-end throughput, and a significant portion can be removed without slowing down training, which we call energy bloat.
@@
In this work, we identify two independent sources of energy bloat in large model training and propose Perseus, a training system that mitigates both. To do this, Perseus obtains the “iteration time–energy” Pareto frontier of any large model training job using an efficient graph cut-based algorithm and schedules the energy consumption of computations across time to remove both types of energy bloat. Evaluation on large models including GPT-3 and Bloom shows that Perseus reduces the energy consumption of large model training by up to 30% without any throughput loss or hardware modification, enabling energy reduction – and therefore cost savings – otherwise unattainable before.
1. Introduction
As deep neural networks (DNNs) continue to grow in model and dataset size (scalinglaws-arxiv20, ; chinchilla-neurips22, ), the energy consumption of large model training is increasing as well. For instance, training GPT-3 (gpt3, ) reportedly consumed 1.3 GWh (patterson2021carbon, ). Then, this was recently dwarfed by a 200B model trained by Amazon consuming about 11.9 GWh (hamilton2024constraint, ), which is enough to power more than 1,000 average US households for a year (us-household, ). Such energy-intensive large model training workloads not only inflate datacenter operational expenses, but also made power delivery a primary challenge in building datacenters today (cbre2023, ; mckinsey2023, ; cbre2024, ).
Despite recent works on accelerating large model training (megatronlm-sc21, ; alpa-osdi22, ; merak-tpds23, ; oobleck-sosp23, ), energy optimization still remains an open challenge (greenai-cacm20, ; patterson2021carbon, ). While energy optimization is well-studied in the hardware community (eyeriss-jssc16, ; vivienne17efficient, ; puma-asplos19, ; kws-jssc24, ), the power bottleneck of recent datacenters (cbre2023, ; mckinsey2023, ; cbre2024, ) shows that efficiency gains from hardware advancement alone are not sufficient to sustain the growing demand for AI compute. As such, a recent line of work demonstrated that software plays a significant role in energy optimization by capturing application characteristics that general-purpose hardware cannot (e.g., no need to finish computation before the deadline), bringing hardware-agnostic energy-efficiency gains (gpoeo, ; zeus-nsdi23, ; chase-ccai23, ; enemy-hpca23, ; envpipe-atc23, ).
@@ 1. Introduction
In this paper, we seek a software method that reduces the energy consumption of large model training without slowdown, thereby also reducing the average power draw. To that end, we identify energy bloat – the portion of energy consumption that can be removed without slowdown – in software systems for large model training. We find two independent sources of energy bloat: intrinsic and extrinsic, and propose a single optimization framework that minimizes both.
@@
Intrinsic energy bloat comes from computation imbalance when a large model is distributed across multiple GPUs with pipeline parallelism (§2.2). Balancing the amount of computation in each pipeline stage is an important problem for distributed execution planning (pipedream-sosp19, ; dapple-ppopp21, ; autopipe-icml21, ; alpa-osdi22, ), but perfectly balancing every stage is not always possible because layers in a DNN are coarse-grained tensor operations with varying amounts of computation. When all stages do not have the same computation time, those not on the critical path of computation run needlessly fast – i.e., they consume energy that does not contribute to the overall training throughput. Such intrinsic energy bloat opens up the opportunity to precisely slow down each non-critical computation in the pipeline such that the length of the critical path does not change.
@@
Extrinsic energy bloat, in contrast, arises when multiple pipelines run in parallel in a synchronous fashion to scale out training to massive datasets, and one or more pipelines run slower than the others (§2.3). Root causes behind such slowdowns are varied, including power/thermal throttling (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-arxiv23, ), I/O bottlenecks in the storage/network (datastalls-vldb21, ; recsys-dsi-isca22, ), hardware/software failures (bamboo-nsdi23, ; oobleck-sosp23, ), etc., and the likelihood of their presence increases with the scale and duration of training (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ). All pipelines running faster than the straggler pipeline are needlessly fast and wasting energy that does not affect the overall training throughput. Thus, extrinsic energy bloat allows us to slow down entire pipelines (while still keeping their intrinsic energy bloat low) without delaying gradient synchronization.
@@
In this work, we propose Perseus, which formulates a unified optimization framework to remove both intrinsic and extrinsic energy bloat from large model training (§3). At its core, Perseus efficiently pre-characterizes the entire “iteration time–energy” Pareto frontier, allowing it to minimize intrinsic bloat under normal operation and to mitigate extrinsic bloat arising from stragglers. Existing works fall short on both fronts. EnvPipe (envpipe-atc23, ) is limited to intrinsic bloat reduction with a point solution that leads to suboptimal energy reduction. Zeus (zeus-nsdi23, ), in contrast, ignores intrinsic bloat as it only considers single-GPU training, which also renders its “iteration time–energy” frontier suboptimal for large models.
@@
We show that characterizing the “iteration time–energy” Pareto frontier is NP-hard not only to solve but also to approximate within any constant factor. Given this impasse, we propose an efficient algorithm that optimally solves a relaxed problem instead (§4). At its core, Perseus represents one training iteration as a directed acyclic graph (DAG). Each node in the DAG represents either a forward or backward computation in a particular stage, and nodes are connected with edges that represent computation dependencies. Perseus’s optimization algorithm then efficiently generates all energy schedules, defined as the planned time and energy consumption of each DAG node (computation), that are on the “iteration time–energy” Pareto frontier. In doing so, instead of finding each point on the frontier individually, Perseus employs a graph cut-based algorithm that iteratively crawls up the frontier from the bottom, generating the next Pareto-optimal energy schedule by carefully tweaking its neighbor energy schedule on the frontier. Minimizing intrinsic and/or extrinsic energy bloat is then as simple as choosing the appropriate energy schedule from the Pareto frontier.
Perseus is split into a server and a client library (§5). The client library can be integrated into any large model training framework as well as any accelerator that has an energy measurement and execution speed control interface, giving the server an abstract view of training computations and their time and energy consumption. The server produces Pareto-optimal energy schedules, which is realized by the client.
@@
Evaluation on large models including GPT-3 (gpt3, ), BERT (bert, ), T5 (t5-jmlr20, ), Bloom (bloom-arxiv22, ), and scaled-up Wide-ResNet (wideresnet-bmvc16, ) shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown, reducing both energy consumption and average power draw (§6).
@@
  • We propose Perseus, a software-only energy optimization system that removes energy bloat through a unified optimization framework and a graph cut-based algorithm.
  • We evaluate Perseus on a diverse set of large model workloads and show that it significantly reduces energy bloat, bringing hardware-agnostic energy savings.
2. Motivation
First, we provide necessary background regarding large model training (§2.1). Then, we introduce intrinsic (§2.2) and extrinsic (§2.3) energy bloat present in large model training, and discuss opportunities for energy reduction (§2.4).
2.1. Large Model Training
Large model training today is mostly dominated by 3D (data, tensor, and pipeline) parallelism (megatronlm-sc21, ; megatron-turing-arxiv22, ; bloom-arxiv22, ; merak-tpds23, ; falcon-arxiv23, ; llama3-blog, ; deepspeed-github, ; megatronlm-github, ; gpt-neox-github, ). Especially, pipeline parallelism partitions a large model and its training batch into multiple stages and microbatches respectively, and pipelines forward and backward computations through stages. Then, such pipelines are replicated multiple times to perform data parallel training. Since all pipelines must synchronize gradients at the end of each iteration, computations for the next iteration can only begin when all pipelines have finished and exchanged gradients.
2.2. Intrinsic Energy Bloat
Figure 1. One training iteration of GPT-3 1.3B with 4 pipeline stages and 6 microbatches on NVIDIA A100 GPUs drawn to scale. For example, F5 and B5 in the S2 row denote Forward and Backward for the fifth microbatch on Stage 2. The critical path of computation is traced with a blue line. Colors show power consumption. Especially, the blue background indicates that GPUs consume power when blocking on communication. Other models are visualized in Appendix A.
Model # Parameters Imbalance Ratio
@@ 2.2. Intrinsic Energy Bloat
Wide-ResNet50 (wideresnet-bmvc16, ) 0.8B 1.23 1.46
Wide-ResNet101 (wideresnet-bmvc16, ) 1.5B 1.09 1.25
@@
Table 1. Latency ratio of the longest to the shortest stage on A100 GPUs. 1.00 would mean perfect balance.
@@
We profile GPT-3 1.3B on NVIDIA A100 GPUs and visualize the timeline of one training iteration in Figure 1(a). In addition to the familiar bubbles in the 1F1B schedule (megatronlm-sc21, ), we observe gaps in between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage. Such gaps exist because the computation time of each pipeline stage is not perfectly balanced. Partitioning stages in a balanced manner is an important problem in distributed execution planning (gpipe-neurips19, ; pipedream-sosp19, ; dapple-ppopp21, ; alpa-osdi22, ), but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying size.
@@
To understand the amount of possible pipeline stage imbalance, we exhaustively searched for the pipeline partition with the smallest imbalance ratio, defined as the ratio of the longest stage forward computation latency to the shortest. [^1] For Transformer-based models, we partition at the granularity of Transformer layers. For Wide-ResNet, we partition at the granularity of bottleneck layers, which are three convolution layers wrapped with a skip connection. Table 1 lists the minimum imbalance ratio for various models, which shows that perfect balance is difficult to achieve. See Appendix B for partitioning details and sources of imbalance.
@@
Given stage imbalance, as we have seen in Figure 1(a), not all forward and backward computations are on the critical path of computation. This means that non-critical computations running at their maximum speed is not contributing to faster iteration time, and thus simply wasting energy. We call this intrinsic energy bloat, which can be reduced by precisely slowing down each non-critical computation without lengthening the critical path (Figure 1(b)). Although seemingly simple, this problem is not only NP-hard to solve, but also NP-hard to even approximate to any constant factor (svensson2012hardness, ).
2.3. Extrinsic Energy Bloat
@@
Numerous replicas of the same pipeline run in a data parallel fashion in large model training. Because every pipeline must synchronize gradients at the end, if a pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes (Figure 2(a)). Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time is wasteful. We call this extrinsic energy bloat, because unlike intrinsic energy bloat, its cause is extrinsic to the pipeline. To reduce extrinsic bloat while keeping intrinsic bloat low, one can determine the energy-optimal iteration time and precisely slow down computations in the non-straggler keeping that iteration time (Figure 2(b)).
@@
Figure 2. Two data parallel pipelines run, and the first one became a straggler (e.g., due to the second GPU temporarily overheating). The non-straggler pipeline wastes energy by running as fast as possible. Such extrinsic energy bloat can be removed by precisely slowing down non-straggler pipelines.
Stragglers arise from numerous sources. Thermal or power throttling in a datacenter can result in 10–50% slowdown (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-arxiv23, ), and I/O bottlenecks in the storage or network can be longer than GPU computation by up to 4×4$\times$4 × (datastalls-vldb21, ; recsys-dsi-isca22, ), acting like a persistent straggler pipeline. Recent failure-resilient training frameworks (bamboo-nsdi23, ; oobleck-sosp23, ) deploy heterogeneous pipelines, introducing non-uniform iteration times. With increasing job and infrastructure scale, the probability of encountering stragglers increases (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ; megascale-nsdi24, ; tpuv4-nsdi24, ).
Stragglers are often times known to and anticipated by the training infrastructure, because they were created by the infrastructure itself (e.g., throttling, non-compute bottlenecks, fault-tolerant planning). Therefore, in this work, we focus on how to plan time and energy across time and allow quick adaptation assuming that such information is available.
2.4. Potential Benefits of Reducing Energy Bloat
To gauge potential energy savings, we measure the energy savings of slowing down every computation in the pipeline to their minimum-energy frequencies. This will slow down the pipeline’s iteration time, but can act as an upper bound for energy savings. For our experimental workloads in Section 6.2, this gives on average 16% and 27% energy reduction on A100 and A40 GPUs, respectively. We show in Section 6.2 that Perseus can realize most of the potential savings with negligible slowdown, even through our problem is NP-hard.
Figure 3. Three cases of where the straggler pipeline’s iteration time $T^{\prime}$ can be. $T_{\min}$ and $T^{*}$ are the shortest and longest iteration time on the “iteration time–energy” Pareto frontier. The black dot is when all computations run at the maximum speed, which wastes energy. The green dot is the energy-optimal iteration time of the non-straggler pipeline. Solid orange dots make up the “iteration time–energy” Pareto frontier, and the orange dotted line shows that energy increases beyond $T^{*}$.
3. Perseus Overview
@@
3.1. Unified Optimization Framework